Use two equations in two variables to solve the application.
A rectangle is 3 times as long as it is wide, and its perimeter is 80 centimeters. Find its dimensions.

Answer by Boreal(15235)

half perimeter= L+W=40
L=3W
so 3W+W=40
W=10 cm
L=30 cm
The perimeter with those two is 2(10)+2(30)=80 cm